OpenStudy (anonymous):

Divide the quantity forty-eight times ten to the eighteenth divided by the quantity four times ten to the fourth. Express the answer in scientific notation.

OpenStudy (anonymous):

@henpen

OpenStudy (anonymous):

Scientific notation problems you want to break it up to make it easier. You have your coefficients and your "10^n"s. Move all your 10^s to the numerator. So rewrite this as: 48/4 * 10^18 * 10^-4 So 48/4 = 12. And 18+(-4) = 14 So your answer is 12 x 10^14

OpenStudy (anonymous):

\[\frac{48*10^{18}}{4*10^4}=\frac{48}{4}*\frac{10^{18}}{10^4}\] Use\[a^b/a^c=a^{b-c}\]